TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Back to overview

YOU ARE VIEWING BlogEntryPoint6x2007x10x09x21x00

From Pressure test to Cool-down - Sector 5-6 _ 09/10/2007

JPEG 5-6.jpg:
5-6.jpg

Feedback of the meeting:

  • Problems with the soldering of the flexible of DFBAJ - To be resolded (after meeting: soldered, waiting for radiography)
  • ElQA on DFBXF.3R5 not yet done (after meeting: possibly done on Thursday)
  • Cool-down could be delayed 1 week (start on w.46)

After meeting:

  • Partial coherence test successfully performed
  • Leak in 15L6 will be repaired on Wednesday; 3 bars will be put during night to check
  • Rail for re-routing of 120 A cables on DFBLE.RR57 is in place; connection on Wednesday

  • Sector 5-6_w41.jpg:
    5-6.jpg
Topic attachments
I Attachment History Action Size Date Who Comment
JPEGjpg 5-6.jpg r1 manage 80.9 K 2007-10-09 - 21:22 UnknownUser  
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-10-09 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login